OpenStudy (anonymous):

Find the intercepts of each line, and graph the line. 5x+6y=30 how do I solve this?

OpenStudy (anonymous):

put x = 0 get y intercept y = 0 get x intercept

OpenStudy (anonymous):

x=0 then 5(0)+6y=30 6y=30 y=30/6 y=5 (0,5) y=0 then 5x+6(0)=30 5x=30 x=30/5 x=6 (6,0)

OpenStudy (anonymous):

@x=0, y=5, @y=0, x=6.

OpenStudy (anonymous):

thank you so much! I have it now thank you!